# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ 5.1 Assumptions
where $S^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$. For any $t>0$, the moment generating generating function of a sub-Gaussian random variable satisfies
$\mathbb{E}[{\sf exp}(t(X_{i}-\mathbb{E}[X_{i}]))]\leq{\sf exp}(\varrho^{2}t^{2}/2)$
where $\varrho$ is the sub-Gaussian constant.
@@
$\mathcal{C}_{k_{0}}(L_{0})=\left\{\nu\in\mathbb{R}^{p}:\|\nu\|_{1}\leq L_{0}\sum_{j=1}^{k_{0}}\left|\nu_{i_{j}}\right|\text{ for some }1\leq i_{1}<\ldots<i_{j}\leq p\right\}$ (5.1)
@@
$\varphi_{X}^{2}(\Sigma_{X},\eta_{0},L_{0}):=\min_{\nu\in\mathbb{R}^{p}}\left\{\frac{\nu^{\top}\Sigma_{X}\nu}{\eta_{0}\|\nu\|_{1}^{2}}:\nu\in\mathbb{R}^{p},\nu\in\mathcal{C}_{k_{0}}(L_{0})\right\}$ (5.2)
@@
We assume that $\beta_{1}$ is $k_{1}$-sparse, i.e., $\|\beta_{1}\|_{0}\leqslant k_{1}$, $\gamma_{1}$ is $k_{2}$-sparse, i.e., $\|\gamma_{1}\|_{0}\leqslant k_{2}$ (which implies $\phi_{1}$ is $k_{1}+k_{2}=k$ sparse), and $\|\hat{\phi}_{1}\|_{0}\leqslant 2(k_{1}+k_{2})$ with high probability [^1] In Step 1 of Algorithm 1, when performing lasso regression of $Y$ on $(X,M)$ on the treated units, we assume that $\hat{\beta}_{1}$ and $\hat{\gamma}_{1}$ are $2k_{1}$-sparse and $2k_{2}$-sparse, respectively, with probability approaching to 1. This implies that $\hat{\phi}_{1}$ is $2(k_{1}+k_{2})$-sparse with high probability. Since the lasso leads to sparsity with high probability, this assumption removes the need for a stronger condition on bounding the $\ell_{1}$-risk of $b-\hat{b}$, which would otherwise require $\mathbf{B}_{0}$ to exhibit column sparsity. . $\mathbf{B}_{0}$ is $s$-row sparse (each row of $\mathbf{B}_{0}$ has at most $s$ active element, i.e., $\|\mathbf{B}_{0,i}\|_{0}\leqslant s$). Furthermore, $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ and $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$.
@@
Assume that the noise is homoskedastic: $\operatorname{Var}[\epsilon_{i}\ |\ \mathbf{W}_{t,i}]=\sigma_{1}^{2}$ for every $i=1,\ldots,n_{t}$, and that the noise $\epsilon_{i}:=Y_{i}^{(1)}-\mathbb{E}[Y_{i}^{(1)}|\mathbf{W}_{t,i}]$ is i.i.d uniformly sub-Gaussian with mean 0 a finite parameter $v_{1}^{2}S_{1}>0$. Similarly, suppose that we have homoskedastic noise: $\operatorname{Var}[U_{i}\mid X_{i}]=\sigma^{2}_{2}$ for all $i=1,\ldots,n_{c}$, and also the response noise $U_{i}$ is i.i.d uniformly sub-Gaussian with 0 mean and a finite parameter $v_{2}^{2}S_{2}>0$.
@@
Consider the random design (whitened) matrices denoted by $\mathbf{Q}=(\mathbf{W}_{t}-\bm{\mu}_{t})\Sigma_{t}^{-1/2}$ and $\mathbf{P}=(\mathbf{X}_{c}-\bm{\mu}_{c})\Sigma_{c}^{-1/2}$ , where $\Sigma_{t}=\operatorname{Var}[\mathbf{W}_{t,i}\mid A_{i}=1]$ is the conditional covariance matrix of $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$ for the treatment observations, $\Sigma_{c}=\operatorname{Var}\left[X_{i}\mid A_{i}=0\right]$ is the conditional covariance matrix for the control observations, and $\mu_{t}=\mathbb{E}[\mathbf{W}_{t,i}\mid A_{i}=1]$ and $\mu_{c}=\mathbb{E}[X_{i}\mid A_{i}=0]$. We denote $\mu=\mathbb{E}[X]$ and $\tilde{a}=\left[\mu\quad\mathbf{B}_{0}\mu\right]^{\top}$. We suppose the following,
  • $\mathbb{E}[Q_{ij}]=0$, $\mathbb{E}[P_{ij}]=0$ and the variance is unity, $\operatorname{Var}[Q_{ij}]=1$, $\operatorname{Var}[P_{ij}]=1$, applicable for every pair of indices $i,j$.
@@
  • Rows of $\mathbf{Q}$ and $\mathbf{P}$, denoted as $Q_{i}$ and $P_{i}$ respectively are all i.i.d sub-Gaussian random vectors with some sub-Gaussian constants $\iota>0$ and $\varsigma>0$.
  • The diagonal of $\Sigma_{t}$ and $\Sigma_{c}$ are entry-wise bounded, hence $(\Sigma_{t})_{jj}\leqslant S_{1}$ for all $j=1,\ldots,p+q$ and that $(\Sigma_{c})_{jj}\leqslant S_{2}$ for all $j=1,\ldots,p$.
  • The eigenvalue of $\Sigma_{t}$ and $\Sigma_{c}$ are lower and upper bounded by some universal constants $(C_{t,\text{min}},C_{t,\text{max}})$ and $(C_{c,\text{min}},C_{c,\text{max}})$, say $\sigma_{\text{min}}(\Sigma_{t})\geq C_{t,\text{min}}>0$, $\sigma_{\text{min}}(\Sigma_{c})\geq C_{c,\text{min}}>0$, and $\sigma_{\text{max}}(\Sigma_{t})\leqslant C_{t,\text{max}}<\infty$, $\sigma_{\text{max}}(\Sigma_{c})\leqslant C_{c,\text{max}}\leq\infty$.
@@
  • The operator norm $\|\mathbf{B}_{0}\|_{\text{op}}\leqslant C$, $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$, $\|\mu\|_{\infty}\leqslant C_{X}$ and $\|\gamma_{1}\|_{2}\leqslant C_{\gamma}$ are finite and upper bounded by some constants $C$, $C_{B}$, $C_{X}$ and $C_{\gamma}$, and $\|\hat{\gamma}_{1}-\gamma_{1}\|_{2}=\sqrt{k\log(q)/n_{t}}$ with high probability.
  • The $\ell_{\infty}$-norm of $(\mu_{t},\mu_{c})$ is bounded, i.e., $\|\mu_{t}\|_{\infty}\leq C_{t}$, $\|\mu_{c}\|_{\infty}\leq C_{c}$ for some constants $C_{t}>0$ and $C_{c}>0$.
@@
  • There exists some constant $\tilde{c}$ and $\kappa$, independent of the sample size $n$ such that $q/p^{\tilde{c}}\rightarrow 0$, and $\tilde{a}^{\top}\Sigma^{-1}_{W}\tilde{a}\leq V_{1}$, $\mu^{\top}\Sigma^{-1}_{c}\mu\leq V_{2}$, i.e., $\tilde{a}^{\top}\Sigma_{t}^{-1}\tilde{a}=\mathcal{O}(1)$, $\mu^{\top}\Sigma_{c}^{-1}\mu=\mathcal{O}(1)$, and the maximum entries of $\tilde{a},\mu$ satisfies $\lVert\tilde{a}\rVert_{\infty},\lVert\mu\rVert_{\infty}\geq\kappa>0$.
@@
Assumption 5.5 and 5.6 are commonly used for analyzing high-dimensional statistical models (e.g., see Bühlmann and Van De Geer, (2011); Lin et al., (2023)). These assumptions impose a restriction on the tail behavior of the errors and the design matrices and facilitate the use of many existing concentration inequalities. In Assumption 5.6, we also assumed that the diagonals of the variance matrices and the operator norm of $\mathbf{B}_{0}$ are finite. The finiteness of the operator norm of $\mathbf{B}_{0}$ implies that the rows of $\mathbf{W}$ are sub-Gaussian, as elaborated in the proof of Theorem 5.11. While it is possible to relax the boundedness assumptions by allowing the bounds to grow with $n$, it would require careful tracking of the constants, as they would affect the estimator’s convergence rate. Since this would not contribute significantly to the core idea of the paper, we choose not to pursue this generalization.
@@
In the setting in our work, it is not reasonable to assume that the coordinates of $\mathbf{W}_{t}$ are independent, as the coordinates of $M_{i}$ depend on each other through $X_{i}$ in our data-generating process. This significantly differs from the theoretical analysis in (Athey et al., , 2018), where the authors assume that individual entries are all independent.
@@
In the proof of the Theorem 5.11, we require a minimum estimand size $\lVert a\rVert_{\infty},\lVert\bar{X}\rVert_{\infty}\geq 2\kappa$. This is required to rule out the fact that if the contrasts are relatively small, then we would end up with having $\tau_{1},\tau_{2}=0$ resulting in a super-efficient estimator as noted in Athey et al., (2018).
@@
We assume that the population covariance matrices $\Sigma_{t}$ and $\Sigma_{c}$ satisfy the $\{k,\eta_{t},L_{t}\}$- and $\{s,\eta_{c},L_{c}\}$-restricted eigenvalue condition respectively, with its corresponding compatibility constant
$\varphi^{2}_{W}(\Sigma_{t},\eta_{t},L_{t})$ and $\varphi_{c}^{2}(\Sigma_{c},\eta_{c},L_{c})$ respectively (see Definition 5.3 for details).
@@ 5.2 Consistency and Asymptotic Normality of $\hat{\theta}_{0}$
$\hat{\theta}_{0}$ $=\bar{X}^{\top}\left(\hat{\beta}_{1}+\hat{b}\right)+\sum_{\{\begin{subarray}{c}i:A_{i}=1\end{subarray}\}}\tau_{1,i}\left(Y_{i}-\mathbf{W}_{t,i}\cdot\hat{\phi}_{1}\right)+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\hat{b}\right)\,,$ (5.3)
@@
with $a=\begin{bmatrix}\bar{X}&(\hat{\mathbf{B}}_{0}\bar{X})\end{bmatrix}^{\top}$, $\hat{\phi}_{1}$ and $\hat{b}$ are sparse linear estimator of $\phi_{1}$ and $b$, respectively, and $K_{1}$ and $K_{2}$ are hyper-parameters (uniformly bounded in $(n,p)$). Equations (5.4) and (5.5) are optimization problems designed to obtain the weights for the bias correction terms in the estimator $\hat{\theta}_{0}$. Thus, it is crucial to establish that these optimization problems are feasible (with high probability), a result we formalize in the following lemma.
@@
Under Assumption 5.6, the optimization problems (5.4) and (5.5) are feasible with probability going 1 for any choice of $K_{1}\geq C\iota^{2}\sqrt{V_{1}^{\prime}S_{1}}$, and $K_{2}\geq C\mathcal{\varsigma}^{2}\sqrt{V_{2}S_{2}}$.
@@
Under Assumptions 5.4, 5.5, 5.6, 5.9, the specification of $(K_{1},K_{2})$ in Lemma 5.10, and the choice $\lambda_{1,n}=\iota^{2}v_{1}\sqrt{(c_{1}^{2}\log(p+q))/n_{t}}$, $\lambda_{2,n}=\varsigma^{2}v_{2}\sqrt{(c_{2}^{2}\log(p))/n_{c}}$, our estimator $\hat{\theta}_{0}$ satisfies:
@@
where $\sigma_{n}^{2}=n\left(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2}\right)$ with $\sigma_{1}^{2},\sigma_{2}^{2}$ are as defined in Assumption 5.5.
@@
When $\gamma_{1}=0$, i.e., the response is unaffected by the mediators, then the Lasso estimator $\hat{\gamma}_{1}$ will be equal to 0 with high probability. In that case, Theorem 5.11 implies $(\hat{\theta}_{0}-\theta_{0})/(\sigma_{1}\|\tau_{1}\|_{2})$ is asymptotically normal, which is Theorem 3 of Athey et al., (2018). Therefore, our result can be viewed as a generalized version of the results in (Athey et al., , 2018) in the presence of high dimensional mediators along with the high dimensional covariates.

### v3

@@ 5.1 Assumptions
where $S^{n-1}$ denotes the unit sphere in $\mathbb{R}^{n}$. See Definition 2.5.6 of Vershynin, (2018) for more details.
@@
$\mathcal{C}_{k_{0}}(L_{0})=\left\{\nu\in\mathbb{R}^{p}:\|\nu\|_{1}\leq L_{0}\sum_{j=1}^{k_{0}}\left|\nu_{i_{j}}\right|\text{ for some }1\leq i_{1}<\ldots<i_{j}\leq p\right\}\,.$ (5.1)
@@
$\varphi_{X}^{2}(\Sigma_{X},\eta_{0},L_{0}):=\min_{\nu\in\mathbb{R}^{p}}\left\{\frac{\nu^{\top}\Sigma_{X}\nu}{\eta_{0}\|\nu\|_{1}^{2}}:\nu\in\mathbb{R}^{p},\nu\in\mathcal{C}_{k_{0}}(L_{0})\right\}\,.$ (5.2)
@@
We assume that $\beta_{1}$ is $k_{1}$-sparse, i.e., $\|\beta_{1}\|_{0}\leqslant k_{1}$, $\gamma_{1}$ is $k_{2}$-sparse, i.e., $\|\gamma_{1}\|_{0}\leqslant k_{2}$ (which implies $\phi_{1}$ is $k_{1}+k_{2}=k$ sparse). $\mathbf{B}_{0}$ is $s$-row sparse ((each row of $\mathbf{B}_{0}$ has at most $s$ active element, i.e., $\|B_{0,i}\|_{0}\leqslant s$)). Furthermore, $(k\vee s)\log(p+q)/\sqrt{n}=o(1)$ and $sk\log(p\vee q)/\sqrt{n_{c}\wedge n_{t}}=o(1)$.
@@
Remark 5.5.
In the proof, we assume that our estimator satisfies $\|\hat{\phi}_{1}\|_{0}\leq 2(k_{1}+k_{2})$ with high probability. To ensure this in practice, an additional step may be applied to threshold out small coordinates of $\hat{\phi}_{1}$; see Zhou, (2010) for more details. The sparsity of $\hat{\phi}_{1}$ helps us to avoid a stronger assumption (column sparsity of $\mathbf{B}_{0}$) for establishing an upper bound on $\|\hat{b}-b\|_{1}$.
@@
We assume that the noise is homoskedastic: $\operatorname{Var}[\epsilon_{i}\ |\ W_{t,i}]=\sigma_{1}^{2}$ for every $i=1,\ldots,n_{t}$, and that the noise $\epsilon_{i}:=Y_{i}^{(1)}-\mathbb{E}[Y_{i}^{(1)}|W_{t,i}]$ is i.i.d uniformly sub-Gaussian with mean zero and $\|\epsilon_{i}\|_{\psi_{2}}\leq v_{1}^{2}S_{1}$. Similarly, suppose that we have homoskedastic noise: $\operatorname{Var}[U_{i}\mid X_{i}]=\sigma^{2}_{2}I$ for all $i=1,\ldots,n_{c}$, and also the response noise $U_{i*}$ is i.i.d uniformly sub-Gaussian with mean zero and $\|U_{i*}\|_{\psi_{2}}\leq v_{2}^{2}S_{2}$.
@@
Consider the random design (whitened) matrices denoted by $\mathbf{Q}=(\mathbf{W}_{t}-\mu_{W,t})\Sigma_{W,t}^{-1/2}$ and $\mathbf{P}=(\mathbf{X}_{c}-\mu_{X,c})\Sigma_{X,c}^{-1/2}$ , where $\Sigma_{W,t}=\operatorname{Var}[W_{t,i}\mid A_{i}=1]$ is the conditional covariance matrix of $\mathbf{W}_{t}=(\mathbf{X}_{t}\quad\mathbf{M}_{t})$ for the treatment observations, $\Sigma_{X,c}=\operatorname{Var}\left[X_{i}\mid A_{i}=0\right]$ is the conditional covariance matrix for the control observations. We suppose the following,
@@
  • $\mathbb{E}[Q_{ij}\mid A=1]=0$, $\mathbb{E}[P_{ij}\mid A=0]=0$ and the variance is unity $\operatorname{Var}[Q_{ij}\mid A=1]=1$, $\operatorname{Var}[P_{ij}\mid A=0]=1$, applicable for every pair of indices $i,j$. In addition, rows of $\mathbf{Q}$ and $\mathbf{P}$, denoted as $Q_{i}$ and $P_{i}$ respectively are all i.i.d sub-Gaussian random vectors with some sub-Gaussian constants $\iota>0$ and $\varsigma>0$.
  • There exists positive constants $s_{1},S_{1},s_{2},S_{2}$ such that $s_{1}\leq\lambda_{\operatorname{min}}(\Sigma_{W,t})\leq\lambda_{\operatorname{max}}(\Sigma_{W,t})\leq S_{1}$ and $s_{2}\leq\lambda_{\operatorname{min}}(\Sigma_{X,c})\leq\lambda_{\operatorname{max}}(\Sigma_{X,c})\leq S_{2}$.
Assumption 5.8 (Model Parameters).
Define $\mu_{W,t}=\mathbb{E}[W_{t,i}\mid A_{i}=1]$, $\mu_{X,c}=\mathbb{E}[X_{i}\mid A_{i}=0]$, $\mu_{X}=\mathbb{E}[X]$ and $a^{*}=\left[\mu_{X}\quad\mathbf{B}_{0}\mu_{X}\right]^{\top}$. We assume the following about the model parameters:
@@
  • $\|\mathbf{B}_{0}\|_{\text{op}}\leqslant C$, $\|\mathbf{B}_{0}\|_{1,\infty}\leq C_{B}$, $\|\mu\|_{\infty}\leqslant C_{X}$ and $\|\gamma_{1}\|_{2}\leqslant C_{\gamma}$ are finite and upper bounded by some constants $C$, $C_{B}$, $C_{X}$ and $C_{\gamma}$. The $\ell_{\infty}$-norm of $(\mu_{W,t},\mu_{X,c})$ is bounded, i.e., $\|\mu_{W,t}\|_{\infty}\leq C_{t}$, $\|\mu_{X,c}\|_{\infty}\leq C_{c}$ for some constants $C_{t}>0$ and $C_{c}>0$, and the maximum entries of $a^{*},\mu_{X}$ satisfies $\lVert a^{*}\rVert_{\infty},\lVert\mu_{X}\rVert_{\infty}\geq\kappa>0$.
@@
  • There exists some constant $\tilde{c}$ and $\kappa$, independent of the sample size $n$ such that $q/p^{\tilde{c}}\rightarrow 0$, and $a^{*\top}\Sigma^{-1}_{W,t}a^{*}\leq V_{1}$, $\mu_{X}^{\top}\Sigma^{-1}_{X,c}\mu_{X}\leq V_{2}$.
@@
Assumption 5.6 and 5.7 are commonly used for analyzing high-dimensional statistical models (e.g., see Bühlmann and Van De Geer, (2011); Lin et al., (2023)). These assumptions impose a restriction on the tail behavior of the errors and the design matrices and facilitate the use of many existing concentration inequalities. In Assumption 5.8, we also assumed that the operator norm of $\mathbf{B}_{0}$ is finite. The finiteness of the operator norm of $\mathbf{B}_{0}$ implies that the rows of $\mathbf{W}$ are sub-Gaussian, as elaborated in the proof of Theorem 5.13. While it is possible to relax the boundedness assumptions by allowing the bounds to grow with $n$, it would require careful tracking of the constants, as they would affect the estimator’s convergence rate. Since this would not contribute significantly to the core idea of the paper, we choose not to pursue this generalization.
@@
In the setting of our work, it is not reasonable to assume that the coordinates of $\mathbf{W}_{t}$ are independent, as the coordinates of $M_{i}$ depend on each other through $X_{i}$ in our data-generating process. This significantly differs from the theoretical analysis in (Athey et al., , 2018), where the authors assume that individual entries are all independent.
@@
In the proof of the Theorem 5.13, we require a minimum estimand size $\lVert a\rVert_{\infty},\lVert\bar{X}\rVert_{\infty}\geq 2\kappa$. This is required to rule out the fact that if the contrasts are relatively small, then we would end up with having $\tau_{1},\tau_{2}=0$, resulting in a super-efficient estimator as noted in Athey et al., (2018).
@@
We assume that the population covariance matrices $\Sigma_{W,t}$ and $\Sigma_{X,c}$ satisfy the $\{k,\eta_{t},L_{t}\}$- and $\{s,\eta_{c},L_{c}\}$-restricted eigenvalue condition respectively, with its corresponding compatibility constant
$\varphi^{2}_{W}(\Sigma_{W,t},\eta_{t},L_{t})$ and $\varphi_{c}^{2}(\Sigma_{X,c},\eta_{c},L_{c})$ respectively (see Definition 5.3 for details).
@@ 5.2 Consistency and Asymptotic Normality of $\hat{\theta}_{0}$
[…]
@@
with $a=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$, $\hat{\phi}_{1}$ and $\hat{b}$ are sparse linear estimator of $\phi_{1}$ and $b$, respectively, and $K_{1}$ and $K_{2}$ are hyper-parameters (uniformly bounded in $(n,p)$). Equations (5.4) and (5.5) are optimization problems designed to obtain the weights for the bias correction terms in the estimator $\hat{\theta}_{0}$. Thus, it is crucial to establish that these optimization problems are feasible (with high probability), a result we formalize in the following lemma.
@@
Under Assumption 5.7 and 5.8, the optimization problems (5.4) and (5.5) are feasible with probability going to 1 for any choice of $K_{1}\geq C\iota^{2}\sqrt{V_{1}^{\prime}S_{1}}$, and $K_{2}\geq C\mathcal{\varsigma}^{2}\sqrt{V_{2}S_{2}}$, where $V_{1}^{\prime}$ and $V_{2}$ are constants defined in Lemma A.1 and Assumption 5.8.
@@
Under Assumptions 5.4, 5.6, 5.7, 5.8, 5.11, the specification of $(K_{1},K_{2})$ in Lemma 5.12, and the choice $\lambda_{1,n}=\iota^{2}v_{1}\sqrt{(c_{1}^{2}\log(p+q))/n_{t}}$, $\lambda_{2,n}=\varsigma^{2}v_{2}\sqrt{(c_{2}^{2}\log(p))/n_{c}}$, our estimator $\hat{\theta}_{0}$ satisfies:
@@
where $\sigma_{n}^{2}=n(\sigma_{1}^{2}\|\tau_{1}\|_{2}^{2}+\sigma_{2}^{2}\|\tau_{2}\|_{2}^{2}\|\gamma_{1}\|_{2}^{2})$ with $\sigma_{1}^{2},\sigma_{2}^{2}$ are as defined in Assumption 5.6.
@@
When $\gamma_{1}=0$, i.e., the response is unaffected by the mediators, then the Lasso estimator $\hat{\gamma}_{1}$ will be equal to 0 with high probability. In that case, Theorem 5.13 implies $(\hat{\theta}_{0}-\theta_{0})/(\sigma_{1}\|\tau_{1}\|_{2})$ is asymptotically normal, which is Theorem 3 of Athey et al., (2018). Therefore, our result can be viewed as a generalized version of the results in (Athey et al., , 2018) in the presence of high-dimensional mediators along with the high-dimensional covariates.